OpenStudy (anonymous):

can someone please help me please ill return the fav Find the slope and y-intercept of the line. y = –2x + 8 (Points : 1) slope: 8, y-intercept: –2 slope: 2, y-intercept: –8 slope: –1, y-intercept: 4 slope: –2, y-intercept: 8 Question 2. 2. Find the slope and y-intercept of the line. y = x – 5 (Points : 1) slope: –5, y-intercept: 1 slope: –1, y-intercept: 5 slope: 5, y-intercept: –1 slope: 1, y-intercept: –5 Question 3. 3. David is selling floral arrangements. Each arrangement uses 1 vase and 12 roses. Each vase cos

OpenStudy (dmndlife24):

y=mx+b where m = slope and b = y-int. Our equation is y = -2x +8 -2 is where m is in the formula....so -2 is the slope. 8 is in the place of the b in the formula so 8 is the y-int.

OpenStudy (anonymous):

i thought that was it but then i looked it up an everyone said it different but now i know thanks sooo much is there any way you could help me with more please

OpenStudy (dmndlife24):

Sure

OpenStudy (anonymous):

aww thanks so much ................................ Find the slope and y-intercept of the line. y = x – 5

OpenStudy (anonymous):

i mean y=x-5

OpenStudy (dmndlife24):

First, can you take a wild guess at what the slope is @vickielt

OpenStudy (anonymous):

is it -5

OpenStudy (dmndlife24):

No, it isn't. However, you have to remember that x does equal 1x....so.... 1 is the slope

OpenStudy (anonymous):

so would the -5 be the yintercept

OpenStudy (dmndlife24):

Exactly :D

OpenStudy (anonymous):

so the slop is 1 and the -5 is the y

OpenStudy (anonymous):

is there anyway you could do that with just 2 more please if you cant i will understand i dont want to bother you to much

OpenStudy (dmndlife24):

Yes, correct. I will say that in the future, you will probably have harder questions such as... -5x + y = 3 In this case you would have to put it in y = mx + b form... y = 5x + 3 And yes, i will try to help you as best as i can.

OpenStudy (anonymous):

sorry my computer was acting crazy but i have had algebra 1 befor an past but my files didt trance over an i forgot lol and thanks youre every kind .............................David is selling floral arrangements. Each arrangement uses 1 vase and 12 roses. Each vase costs David $2.00. Let C be the total cost of the arrangement and r be the cost of 1 rose. Which equation should David use to find the total cost of each arrangement?

OpenStudy (anonymous):

an i am bout to ost another question in one sec i might take a few sec to responed Transform this equation to slope-intercept form: 6x + 4y = 8

OpenStudy (dmndlife24):

@vickielt And do you think you could post these two questions in a new question :)

OpenStudy (anonymous):

what d you mean i dont understand

OpenStudy (dmndlife24):

Simply post the question above as a new question

OpenStudy (anonymous):

oh okay on sec please

OpenStudy (anonymous):

David is selling floral arrangements. Each arrangement uses 1 vase and 12 roses. Each vase costs David $2.00. Let C be the total cost of the arrangement and r be the cost of 1 rose. Which equation should David use to find the total cost of each arrangement

OpenStudy (anonymous):

Transform this equation to slope-intercept form: 6x + 4y = 8.

OpenStudy (anonymous):

like this ??

OpenStudy (dmndlife24):

Do you see the "Ask a question..." button above all the other questions

OpenStudy (anonymous):

do you mean make a new post ??

OpenStudy (dmndlife24):

Yeah

OpenStudy (anonymous):

oh okay one sec lol